The difference of two numbers is 18. The sum of the numbers is 84. What is the larger number? What is the smaller number?
Answer by macston(5194)

x-y=18 eq1
x+y=84 eq2
Solve eq1 for x
x-y=18 add y to both sides
x=18+y
Substitute for x in eq2
18+y+y=84 subtract 18 from each side
18+y+y-18=84-18
y+y=66
2y=66 divide both sides by 2
y=33 ANSWER 2 Smallest Number
x-y=18
x-33=18
x=51 ANSWER 1 Largest number
CHECK:
x+y=84
51+33=84
84=84